### v2

# Exploring the spectral characteristics of the periodic burster 4U 1323-62: Type-I X-ray burst and persistent emission

Abstract
We report on the results obtained by the analysis of persistent and type-I thermonuclear X-ray burst emission observed from the periodic burster 4U 1323-62. These analyses are based on the NuSTAR observation performed on 2024 August 7 for a total exposure of around 90 ks. The persistent emission is well described by an absorbed thermal Comptonization model. An absorption edge is also detected at an energy of 7.87 keV, which indicates the presence of absorbing material in the vicinity of this system. Six bursts have been observed during this observation, wherein we find the burst recurrence time to be hr. All the bursts exhibit the characteristics of a sharp rise and exponential decay. We perform the time-resolved spectroscopy of the burst spectra described by a model consisting of thermal emission from the neutron star surface and a varying persistent emission component to study the evolution of burst parameters. The enhancement of the persistent emission during burst exposure is characterized by the scaling parameter , which reflects the increasing strength of the burst–disc interaction with burst intensity, likely driven by Poynting–Robertson drag. The spectral analysis of bursts estimate the average apparent blackbody emitting radius of the neutron star to lie within 1.5–3.5 km. The ignition depths computed from the burst parameters indicate short Type-I thermonuclear bursts from a mixed hydrogen-helium fuel layer.
Keywords: accretion - Neutron star: X-ray Binary – Neutron Star Low Mass X-ray binary – X-ray bursts: spectral analysis – Periodic burster: timing analysis – individual 4U 1323-62
1 Introduction
In a low-mass X-ray binary (LMXB), the neutron star accretes matter from a low-mass companion () via Roche lobe overflow, forming an accretion disk where viscous dissipation heats the inflowing gas (1987huba.conf...35B). As material accumulates on the stellar surface, unstable nuclear burning ignites once the accreted mass exceeds a critical limit, producing a thermonuclear runaway (2021ASSL..461..209G). Type-I thermonuclear bursts occur when the X-ray luminosity of a neutron star rises abruptly and decays over a longer timescale (2017A&A...606A.130I). The burst power generally increases rapidly from the persistent emission level (1036 erg s-1) to about 1038 erg s-1 within a few seconds, and subsequently declines exponentially over a timescale of seconds to minutes—exhibiting the typical fast-rise, exponential-decay (FRED) profile (2017MNRAS.468.4735L).
The accreted hydrogen- and helium-rich material becomes compressed into a dense, thin layer under the neutron star’s intense gravitational field (BERTULANI201656). Continued accretion raises the temperature and density with increasing depth, aided by heat conduction from the stellar core. As matter accumulates, the base of this layer eventually reaches an ‘ignition depth’ where the temperature and pressure become sufficient to trigger thermonuclear fusion. The resulting energy generation—primarily from the triple-alpha process and the hot CNO cycle—further increases the local temperature, thereby accelerating the fusion rate (kuulkers2003photospheric).
During a burst, a huge amount of energy (1039 erg) is released almost instantaneously, causing the heated photosphere to cool over time and producing the characteristic sharp rise and gradual decay in X-ray flux (WIESCHER200751). The interval between successive bursts, referred to as the recurrence time, is closely associated with the parameter , defined as the ratio of persistent (accretion) to burst emission flux (1993ApJ...413..324T; wang2024burst). When the energy released by thermonuclear runaway exceeds the neutron star’s gravitational binding energy, radiation pressure drives the photosphere outward, resulting in photospheric radius expansion (PRE). This occurs when the burst luminosity temporarily reaches or surpasses the Eddington limit, altering the local mass accretion rate. As the radiation pressure declines, the accreting matter experiences Poynting–Robertson drag, modifying both the accretion flow and luminosity (1977cami.coll..121S). Subsequently, the radiation energy dissipates, the photosphere cools and contracts back to the stellar surface (10.1093/mnras/220.2.339). In the absence of huge radiation pressure, the gravitational pull restores the accretion rate to its pre-burst steady level (Peng_2025).
The source 4U 1323-62 was first detected by Uhuru and Ariel V missions in 1970s which were later reported in 1978 and 1981, respectively (forman1978fourth; warwick1981ariel). The source was classified as a neutron star on the first detection of a burst using EXOSAT and one hour intensity dips were reported in 1985 (1985SSRv...40..287V). One X-ray burst was reported during the dip in the 2-10 keV energy regime with peak flux 7 10-10 ergs cm-2 s-1. The source is also known as XB 1323-619 and it has been reported to show intensity dips and X-ray bursts (balucinskachurch1999xraystudydippinglow; barnard2001; boirin2005highly; balucinska2009neutral; Bhulla_2020). During the BeppoSAX observation of the source in 1997, balucinskachurch1999xraystudydippinglow reported a total of 10 X-ray bursts and 12 intensity dips. Apart from 4U 1826-24, it is the other source that is designated as a ’clocked’ burster as their bursting remains periodic over extended periods of time (balucinska2009neutral). The dipping occurs due to absorption of the emitted X-ray by the orbital obscuration of the X-ray emitting region by an absorbing region which is structured azimuthally around the neutron star (1982ApJ...253L..61W). The dipping periodicity indicates that the source is viewed nearly in edge-on orientation (), hence, the inclination angle (angle between the direction of the line-of-sight and rotation axis of the accretion disk) (1987A&A...178..137F). barnard2001 reported on the 1997 observation of both BeppoSAX and RXTE. For RXTE observation, they reported seven bursts and seven dips that repeated almost periodically. The last AstroSat observation of 4U 1323-62, during the period 2016-17, reported six bursts with recurrence time 9400 s and two dips with time separation 11.928 hr (Bhulla_2020). The irregular intensity dips were previously reported with a recurrence time 2.94 hr from the INTEGRAL observation with no variation in the dips within 20-40 keV energy regime (balucinskachurch1999xraystudydippinglow). boirin2005highly reported presence of Fe xxv and Fe xxvi 1s-2p absorption lines around 6.9 keV which are similar to the values reported from XMM-Newton observation by Church_2005.
Non-dip spectra was modelled via two components, one of which was absorbed and the other was unabsorbed, it is known as ’absorbedunabsorbed’ approach (1986ApJ...308..199P; smale1992ginga). In the 1985 observation by 1985SSRv...40..287V, the blackbody component for the burst spectrum was fitted that showed cooling during the burst decay. It was reported to have an overall flux of 7-8 10-11 ergs cm-2 s-1 in the energy range 2-10 keV with a power law index ranging between . 1997ApJ...491..388C showed that in order to fit the spectral evolution for the emission after dipping, the resultant model requires emission of two components to fit the dip states as well as the spectra formed at the non-dipping states. It is shown previously that the best-fit model, called the ’complex continuum’ model, consists of a component to fit the Comptonized emission from the accretion disk corona (extended region) and a blackbody component to fit the emission from the neutron star (point-source) (balucinskachurch1999xraystudydippinglow; barnard2001). balucinskachurch1999xraystudydippinglow reported that for the broadband study of the source in the 0.1-150 keV energy regime, the non-bursting non-dipping intervals could be modelled with a power law with a photon index =, a cut-off energy keV. And, the fit for the blackbody emission provided with the blackbody temperature k keV. The fitting of the persistent emission showed that in the 2-10 keV energy range, the blackbody contribution in the total flux is 15% whereas during dips, the maximum reduction in intensity is 65%. barnard2001 reported the blackbody temperature for the RXTE observation as k keV and the cut-off power law index, . In the RXTE observation out of the seven bursts, two bursts were reported to be double burst. Later, BOIRIN20062759 proposed a photo-ionized absorber model rather than individual Gaussian profile to explain the spectral evolution from persistent to dipping LMXB. Frequency-resolved analyses of the source have previously revealed a low-frequency QPO at 1 Hz (jonker1998persistent; balman2010; Bhulla_2020).
In the present work, we have carried out a detailed persistent spectral analysis and time-resolved spectral analysis of the short successive type-I X bursts observed during the NuSTAR observation. This paper is structured as follows. In section 2, we state the observation details and describe data reduction process. Throughout section 3 we discuss the light curves for the observation for the observed bursts and intensity dips. In section 4 we present the adapted models for the persistent and burst emission to carry out the spectral analysis. Additionally, we discuss the time-resolved analysis of three bursts. Section 5 discusses the observations and inferences drawn from the spectral analysis.
2 Observation and Data reduction
The analysis presented in this work is based on observations of 4U 1323-62 conducted by NuSTAR on 2024 August 07 (MJD ) for an effective exposure of ks under the archival Observation ID: 31001021002. The Nuclear Spectroscopic Telescopic Array (NuSTAR), launched by NASA on 2012 June 13, is the first focusing hard X-ray telescope operating in the high energy ( keV) range (harrison2013nuclear).
The NuSTAR data were processed using the data analysis software NuSTARDAS v2.1.5, within HEASOFT v6.35.1 employing the latest calibration database (CALDB v20250415). The task nupipeline v0.4.12 was used to generate the calibrated and screened event files. A circular extraction region of 120” radius was used to study the persistent and burst spectrum. From the same chip, the background was selected as a region of radius 120” far away from the source. HEASOFT provides FTOOLS to deduce the FITS files needed for analysing the observed data. FTOOL nuproducts was used to extract the spectra and light curves from FPMA and FPMB. For time-resolved spectroscopy, good time interval (GTI) files were created separately for the persistent emission and the observed type-I X-ray bursts. During burst analysis, data above 20 keV were excluded due to background dominance.
3 Light Curve
The complete NuSTAR FPMA light curve in the 3–79 keV band is shown in Fig. 1. The average count rate during the persistent phase was 3-5 counts s-1. Overall six type-I thermonuclear bursts were detected with their timing properties listed in Table 1. Alongwith the six bursts, seven regular intensity dips have been observed.
We also plotted two light curves with 100 s bin size in the 3–79 keV energy band, using time intervals of the first 27 ks and from 40 ks to 80 ks, which spans a total of 40 ks. Those are presented in the right panel of Fig. 1. From the light curves, we found evidence of regular intensity dips with a period of approximately 2.91–2.98 hours (indicated by green double-headed arrows). The source is known for periodic intensity dips, with a period of 2.94 hours, first discovered through EXOSAT observations and later confirmed by other observations (1985SSRv...40..287V; boirin2005highly; Bhulla_2020).
The X-ray bursts as shown in Fig. 2 have the characteristic of a fast rise and exponential decay (FRED). The peak count rate during bursts B1, B2, B3, B4, B5 and B6 (in unit of counts s-1) are 144, 178, 196, 94, 146 and 154 respectively. The peak count rates are at an average 25-50 times higher than the persistent level. In Fig. 2 for each occasion of burst, we have considered the initiation point of the burst where the count rate is 25% of the peak count rate of that particular burst (shown by the green dotted-line). Rise time for the burst is considered as the time taken to increase from the initiation level to the level of 90% of the peak count rate. The time duration for peak of the burst is considered as the total time for which the count rate remains above 90% of the peak count rate following galloway2008thermonuclear. Decay time for the burst is considered to be the time taken for the count rate to fall down below the initiation level. The total integrated time for each of the bursts is summation of its rise time, peak time and decay time. We have obtained the integrated times for B1, B2, B3, B4, B5 and B6 to be 105 s, 107 s, 109 s, 36 s, 108 s and 91 s, respectively (Table 1).
The rise times of bursts B1–B6 were measured to be 7 s, 4 s, 7 s, 2 s, 7 s, and 4 s, respectively, while their decay tails were significantly longer—95 s, 102 s, 101 s, 33 s, 99 s, and 83 s. To examine their spectral evolution, we conducted time-resolved spectral analysis for all bursts. Each of the bursts, barring B4 and B6, have exposure above 100 s. For B4 and B6, the burst exposures were shorter (36s and 91s, respectively). A time segment of 20 ks duration, from elapsed time 30th ks to 50th ks, was used to ascertain the flux parameter of persistent emission.
| Burst Serial no. | Start time (s) - End time (s) | trise (s) | tpeak (s) | tdecay (s) | Burst exposure (s) | Peak intensity (counts s-1) |
| B1 | 6430 - 6550 | 7 | 3 | 95 | 105 | 144 |
| B2 | 24000 - 24120 | 4 | 1 | 102 | 107 | 178 |
| B3 | 53590 - 53710 | 7 | 1 | 101 | 109 | 196 |
| B4 | 69570 - 69620 | 2 | 1 | 33 | 36 | 94 |
| B5 | 101380 - 101500 | 7 | 2 | 99 | 108 | 146 |
| B6 | 116520 - 116630 | 4 | 4 | 83 | 91 | 154 |
4 Spectral Analysis
The spectral analysis for the persistent and burst emission was carried out using XSPEC v12.13.0c (1996ASPC..101...17A). We carried out the time-resolved analysis for three bursts (B1, B2 and B5) having five segments for each in the energy range 3-20 keV. To account for the interstellar absorption along the line of sight, we have used the TBabs model. During the analysis, the neutral hydrogen column depth density is kept constant at (1985SSRv...40..287V; 1990ARA&A..28..215D). The source distance is considered to be 10 kpc away (Bhulla_2020; 1995AJ....110.1674C; 1996rftu.proc..141G).
4.1 Persistent spectral analysis
We extracted the persistent spectrum of the source from a 20 ks time interval (from the elapsed time of 30th ks to 50th ks) of the current observation using the NuSTAR FPMA and FPMB data. No X-ray bursts and prominent intensity dips have been observed in this time interval. We fit them simultaneously in the energy band 3 to 60 keV, where the source counts are well above the background counts. We include a constant multiplication factor to account for cross-calibration of the two instruments, FPMA and FPMB.
Initially, we attempted to fit the persistent spectrum with an absorbed cutoff power law model (Tbabs*cutoffpl), which is referred to as Model 1. The Model 1 resulted in = 599/518 with a photon index of = 1.50 0.02 and the cutoff energy Ecut = 80 keV. Previous studies of this source also reported that the energy spectrum is dominated by a power-law component. The measured parameter values, and Ecut, are consistent with values obtained from BeppoSAX and Suzaku observations (boirin2005highly; balucinska2009neutral).
To describe the persistent emission more precisely, we employed a thermal Comptonization model, nthcomp (1996MNRAS.283..193Z; 1999MNRAS.309..561Z), which involves a seed photon source as a blackbody or a disk-like multi-color blackbody. These seed photons are up-scattered by hot electrons, resulting in the output spectrum. We assumed that the source of the seed photons is the accretion disk ( inptype = 1). This model Tbabs*nthcomp (Model 2) resulted in = 550/517 ( = -49 for 1 additional parameter, as compared to Model 1) with a photon index of , the electron temperature kTe=31 keV, and the seed temperature kTbb = 1.31 0.06 keV. The electron temperature of the corona is comparable to the observed cut-off energy of the spectrum. In addition, we found an absorption feature at 7-8 keV and tried to fit it with an absorption edge model. The inclusion of the edge model, Tbabs*edge*nthcomp (Model 3), improved the fit significantly to = 511/515 ( = -39 for 2 additional parameters, compared to Model 2) with edge energy E = 7.87 0.11 keV and optical depth = 0.12 0.02.
To provide a more accurate model description, we further replaced nthcomp with the convolution model thcomp (1996MNRAS.283..193Z; 2019MNRAS.485.2942N; 2020MNRAS.492.5234Z), an updated version of nthcomp that aligns more closely with the actual Monte Carlo spectra from Comptonization than nthcomp does. The model thcomp involves the upscattering fraction (fsc) of photons from any seed emission and gives the low-energy cut-off and the high-energy rollover corresponding to the soft tail of the seed photon and electron temperature of the corona, respectively. As we assumed that the accretion disk supplies the seed photons, we applied the model Tbabs*edge*thcomp*diskbb (Model 4). This model also satisfactorily described the persistent spectrum, with = 506/514, and provided similar best-fit parameter values, including absorption edge energies compared to Model 3, suggesting that the source geometry is correctly identified. All the best-fit parameters for the persistent spectral fitting are listed in Table 2. The persistent spectra for the best fit values corresponding to Model 3 is shown in Fig. 3.
We also explored the pre-burst persistent spectra of the source before examining the X-ray burst emissions in detail. For this, we extracted the spectra of the persistent emission from exposures of 600, 600, 1000, 300, 1000, and 1000 s, respectively, just before the onset of bursts B1, B2, B3, B4, B5, and B6, depending on the data availability on the same NuSTAR orbit. The 3 – 60 keV energy spectrum of the pre-burst emission was also successfully fitted with an absorbed thermal Comptonization model along with an absorption edge (Model 3). We note that the best-fit parameter values are similar within error when comparing the values obtained from the persistent emission of the 20-ks exposures described earlier. However, parameters such as electron temperature (k) and edge energy (Eedge) are not well constrained by the pre-burst persistent spectra due to the limited exposures and low statistics (low count rate) before X-ray bursts. We fixed them to the best-fit values obtained earlier. The power-law photon index , seed photon temperature k, absorption optical depth , and the corresponding values obtained from modeling the pre-burst emissions are reported in Table 3.
| Model no. | Component | Parameter (unit) | Persistent |
| 1 | constant | FPMB (wrt FPMA) | |
| Tbabs | (f) | ||
| cutoffpl | |||
| Norm () | |||
| / | 599/518 | ||
| 2 | nthcomp | ||
| / | 550/517 | ||
| 3 | edge | E | |
| Max | |||
| nthcomp | |||
| Norm | |||
| / | 511/515 | ||
| 4 | thcomp | ||
| Covering fraction (fsc) | |||
| diskbb | |||
| Norm | |||
| / | 506/514 | ||
| Note: | refers to the degrees of freedom ascertained by | ||
| the components for the particular best-fit model. | |||
| Before Burst | k | |||
| 1 | 1.74 0.05 | 1.51 0.31 | 0.17 0.09 | 45 / 57 |
| 2 | 1.860.5 | 43 / 49 | ||
| 3 | 1.74 0.04 | 1.59 0.20 | 0.19 0.07 | 104 / 90 |
| 4 | 1.77 0.07 | 31 / 25 | ||
| 5 | 1.72 0.03 | 1.35 0.25 | 0.11 0.05 | 86 / 98 |
| 6 | 1.76 0.05 | 1.79 0.17 | 0.14 0.07 | 90 / 90 |
4.2 Burst spectral analysis
Initially, we extracted the spectra for all the individual X-ray bursts (Table 1) and performed simultaneous spectral fitting in the energy range of 3–20 keV. We used the absorbed blackbody model, Tbabs*bbodyrad (Model 5), to fit the burst spectra. The bbodyrad model has two parameters, the blackbody temperature, kTbb, and the normalization. We found that the blackbody Model 5 yielded a poor fit of = 646/458. For B1, Model 5 failed to reproduce the burst spectra at both low as well as high energy ranges and for the other bursts, discrepancies appeared mainly at higher energies. To address the observed excesses, we incorporated the thermal Comptonization component nthcomp along with the edge component resulting in composite model Tbabs*(edge*nthcomp+bbodyrad) (Model 6). All the parameters of edge and nthcomp were fixed to their best-fit values corresponding to persistent emission (Table 2), excluding the parameter norm of nthcomp, which was allowed to vary during spectral fit. Model 6 significantly improved the fit, resulting in a of 530/457 ( for 1 additional parameter, compared to Model 5).
The burst spectra are shown in Fig. 4. As a best-fit value for all the bursts from Model 6, we obtained the blackbody temperature as k = 1.802 0.027 keV and radius for the blackbody emitting region km. The nthcomp norm is obtained as . The time-resolved spectral analysis was performed following a similar approach to that adopted by mondal2025nustar for the “clocked burster” GS 1826–24.
We tested other scenarios to fit the burst spectra apart from the model. First, we attempted to fit the burst spectra by the model TBabs*thcomp*bbodyrad, considering that the burst photons could be affected by the corona/boundary layer (2022ApJ...936...46C). Here, thcomp (2020MNRAS.492.5234Z) is a more accurate version of nthcomp, and the parameter values of this model are extracted from the persistent fit. However, this model failed to fit the residuals observed at low and high energies and provided a fit to the burst spectra that was significantly worse compared to the previous model. Second, we tried the model TBabs*(bbodyrad+powerlaw+nthcomp) to fit the burst spectrum, in which the nthcomp parameter values were fixed to the persistent emission (2025ApJ...986...16J). However, the current data failed to constrain all the free parameters simultaneously and could not produce conclusive results. Third, to fit the burst spectra, another combination of models like TBabs*(bbodyrad+
relxillNS+nthcomp) can be tested in which relxillNS accounts for the reflection of the burst photons from the accretion disk (2023A&A...670A..87L). However, we could not implement this model because of the short exposure time of each segment of the burst spectrum and the lack of a priori knowledge of some important disk-related parameters. The above-mentioned facts restricted us from discussing these models.
4.3 Time-resolved analysis of the burst spectra
The burst exposures for B1–B6 were divided into shorter time segments to track the evolution of spectral parameters. Except for B4 and B6, each burst lasted over 100 s and was divided into five 20 s segments, with any remaining time beyond 100 s included in the final segment (S5). The segments were labeled sequentially as S1–S5. Due to limited data points and poor spectral fits, B4 and B6 were excluded from the time-resolved analysis. Additionally, B3 was omitted because its later segments (S3–S6) had significantly fewer counts compared to the corresponding intervals in B1, B2, and B5.
We did time-resolved spectroscopy for three of the observed six bursts, namely B1, B2 and B5. Initially, we used the model Tbabs*bbodyrad (Model 5) to fit the burst spectra. The corresponding best-fit spectra for B1, B2 and B5 are presented in the upper panels of Fig. 5, 6 and 7, respectively. It is evident that the single-component bbodyrad model does not adequately reproduce the observed spectra, as significant residuals are present in the higher energies. Following the approach of worpel2013evidence, the burst spectra were subsequently fitted with a two component model comprising a blackbody component bbodyrad to describe the burst emission and the persistent emission is modeled using edge*nthcomp, scaled by a variable multiplicative factor () that represents the enhancement of the persistent emission. The parameters for bbodyrad component and were allowed to vary freely. We used the model Tbabs*(*edge*nthcomp+bbodyrad) (Model 7) to analyse the burst emission. A value of indicates that the persistent emission during the burst is identical to its pre-burst level; while implies an enhancement of the persistent emission during burst. The nthcomp parameters redshift and inp_type were set to 0 and 1, respectively, while all other parameters of the edge and nthcomp components were fixed to their best-fit values obtained from the persistent emission fit (see Table 3). The unabsorbed fluxes for the total, Comptonized, and blackbody components were derived using the XSPEC convolution model cflux applied to the relevant spectral components.
| Component | Parameter (unit) | S1 (20s) | S2 (20s) | S3 (20s) | S4 (20s) | S5 (40s) |
| tbabs | ∗ | 1.2 (f) | 1.2 (f) | 1.2 (f) | 1.2 (f) | 1.2 (f) |
mult.
factor |
| cflux | ||||||
| bbodyrad | (keV) | |||||
| Norm | ||||||
| cflux | ||||||
| cflux | ||||||
| 4.89 / 8 | 46.23 / 41 | 18.15 / 18 | 4.76 / 9 | 14.48 / 12 | ||
| Note: | † denotes the unit of flux in erg cm -2 s -1. | |||||
| Component | Parameter (unit) | S1 (20s) | S2 (20s) | S3 (20s) | S4 (20s) | S5 (40s) |
| tbabs | ∗ | 1.2 (f) | 1.2 (f) | 1.2 (f) | 1.2 (f) | 1.2 (f) |
mult.
factor |
| cflux | ||||||
| bbodyrad | (keV) | |||||
| Norm | ||||||
| cflux | ||||||
| cflux | ||||||
| 13.21 / 12 | 33.44 /41 | 15.43 / 18 | 10.95 / 11 | 11.83 / 17 | ||
| Note: | † denotes the unit of flux in erg cm -2 s -1. | |||||
The best-fit parameters of the B1 are shown in Table 4, and the corresponding spectra are shown in lower panel of Fig. 5. The blackbody emitting radius of the neutron star reaches a maximum 3.67 km during S2, where the bbodyrad flux peaks at ergs cm -2 s -1 and then gradually decreases. However, for some segments we obtained a low reduced value because of insufficient count rates.The blackbody temperature (k) for B1 reaches a maximum at S2 2.04 keV and then decreases monotonically till S5. The best-fit parameters for time-resolved spectroscopy of B2 are listed in Table 5 and the corresponding spectra are shown in the lower panel of Fig. 6. The bbodyrad flux reaches its peak value during segment S2, after which the bbodyrad flux steadily decreases through S5. The apparent blackbody emitting radius of the neutron star is estimated around 3.09 km during S2, increasing slightly to 3.29 km in S3. A larger radius of 3.71 km is obtained for S5; however, this value is poorly constrained due to limited statistics (0.696), and thus is not considered reliable. The evolution of the blackbody temperature () shows a maximum of keV in S2, followed by a gradual decrease toward S5. The best-fit spectral parameters for B5 across its five time segments are presented in Table 6, with the corresponding spectra and residuals shown in the lower panel of Fig. 7. The blackbody flux reaches its maximum during segment S2 ( erg cm -2 s -1) and subsequently declines in the later segments. The evolution of the blackbody radius closely follows the flux variation, attaining its largest value of approximately 3.5 km during S2, while the blackbody temperature () also attains its highest value at this segment. Across all bursts, the persistent emission shows a pronounced enhancement (), ranging within 1–4 in B1, 1–8 in B2, and 1–7 in B3, suggesting a variable mass accretion rate during the burst evolution (worpel2013evidence).
| Component | Parameter (unit) | S1 (20s) | S2 (20s) | S3 (20s) | S4 (20s) | S5 (40s) |
| tbabs | ∗ | 1.2 (f) | 1.2 (f) | 1.2 (f) | 1.2 (f) | 1.2 (f) |
mult.
factor |
| cflux | ||||||
| bbodyrad | (keV) | |||||
| Norm | ||||||
| cflux | ||||||
| cflux | ||||||
| 23.10 / 19 | 36.55 / 36 | 7.26 / 15 | 8.39 / 10 | 12.80 / 10 | ||
| Note: | † denotes the unit of flux in erg cm -2 s -1. | |||||
5 Discussion
We report on the NuSTAR observation of the periodic burster 4U 1323-62, performed on 2024 August 7, for a total exposure of 90 ks. We observed the presence of six thermonuclear X-ray bursts and regular intensity dips in the NuSTAR light curve of this source. The observed burst profiles are very similar to those extracted earlier from the XMM-Newton EPIC PN and AstroSat LAXPC detectors during a persistent interval (boirin2005highly; Bhulla_2020). The observed period ( 2.91-2.98 hr) of the intensity dips is also consistent with the other observations (1985SSRv...40..287V; boirin2005highly; Bhulla_2020). During this observation, the persistent emission is well described by the absorbed thermal Comptonization model in the energy band of 3–60 keV. Our persistent spectral analysis suggests that the Comptonization of soft photons originating from the accretion disc is occurring. The power law index , electron temperature of the corona k, and the seed photon temperatures kTbb obtained from the Comptonization model are 1.72 0.01, 23 4 keV, and keV, respectively. In addition, we detected an absorption edge at an energy of 7.87 keV, indicating the presence of an absorbing medium in the vicinity of the system. Apart from the dip intervals, the persistent flux remains unchanged on our observed timescale. We measured a 3-60 keV persistent flux of F erg cm-2 s-1, which corresponds to a persistent luminosity of Lper 3.321036 erg s-1 for an assumed distance of 10 kpc.
Of the six bursts, we performed time-resolved spectral analysis for three- B1, B2, and B5-while the remaining bursts lacked sufficient data points or yielded poor fit values (/dof). Each time-resolved spectrum was fitted with Model 7, which incorporates both the burst and enhanced persistent emissions. The evolution of blackbody temperature, flux, normalization, and the scaling factor () for these bursts is shown in Fig. 8.
In all three cases, the maximum blackbody flux occurs during segment S2, corresponding to the burst peak. The maximum peak fluxes (in units of erg cm-2 s-1) are for B1, for B2 and for B5 indicating a gradual decline from B1 to B5. The normalization parameter (norm), which traces the apparent radius of the blackbody-emitting region, peaks at S2 for both B1 and B5, decreases to a minimum at S4, and increases slightly at S5 (Fig. 8). However, the norm values at S5 for each burst carry large uncertainties—e.g., for B2 and for B5 (Table 5)—indicating limited reliability. The norm values at S2 () and S3 () in B2 correspond to blackbody radius of 3.09-3.29 km. The similarity of S2 and S3 values suggests that the norm for B2 shows a similar evolution pattern to B1 and B5. The maximum estimated blackbody radii are km for B1, km for B2, and km for B5.
The evolution of the blackbody temperature (k) for bursts B1, B2, and B5 (Fig. 8) shows a rise during the burst phase followed by a decline during decay, with the maximum temperature occurring in segment S2. Since the burst rise (S1) transitions into the peak at the S1-S2 boundary, the k values for these two segments are nearly identical. The peak temperatures are keV, keV, and keV for B1, B2, and B5, respectively (Tables 5, 6, and 7). The persistent flux () exceed the blackbody flux while approaching towards the final segment (S5) of the burst exposure as indicated by the best-fit spectral parameters for B1, B2 and B5 corresponding to Model 7 (see the lower panels of S5 in Fig. 5, 6, 7). This growing dominance of the persistent emission implies enhanced up-scattering of low-energy photons, consistent with the mechanism described by 1966ApJ...145..560W.
The scaling factor () quantifies the enhancement of persistent emission relative to its pre-burst level. From the time-resolved spectral studies of the bursts with the model, we found a significant enhancement of the persistent emission during the burst, quantified by . The strength of the persistent emission enhancement scales directly with the burst luminosity. For all three bursts, reaches its maximum in segment S2, coinciding with the burst peak. The peak values are for B1, for B2, and for B5. For B2, the most powerful burst amongst B1, B2, and B5 required a substantial enhancement () of the persistent emission during the burst peak. This trend provides quantitative evidence for a physical scenario where the strength of the burst-disc interaction, likely caused by the Poynting-Robertson drag, is a direct function of the thermonuclear flash’s radiative power (worpel2013evidence; Worpel_2015).
| Burst Serial no. | Bolometric flux† during the burst in the energy range keV | Burst fluence | Decay time of the burst (s) | ergs | Ignition depth ()% | ||||
| B1 | 13.95 | 1.46 | 31.44 | 46.58 | 17.52 | 3.36 | 0.63 | 62.31 | 2.99 |
| B2 | 13.61 | 1.46 | 38.89 | 37.43 | 17.42 | 3.03 | 0.56 | 56.57 | 3.30 |
| B3 | 14.43 | 1.57 | 30.93 | 50.84 | 18.82 | 3.13 | 0.58 | 53.98 | 3.45 |
| B4 | 10.70 | 0.39 | 13.05 | 29.52 | 4.61 | 3.09 | 0.58 | 217.52 | 0.86 |
| B5 | 13.29 | 1.44 | 32.73 | 43.85 | 17.17 | 3.35 | 0.63 | 63.44 | 2.94 |
| B6 | 12.91 | 1.18 | 30.35 | 38.72 | 14.06 | 3.22 | 0.60 | 74.46 | 2.50 |
| Note: | † denotes the unit of flux erg cm -2 s -1; | $ denotes the unit in erg cm-2; | % denotes the unit of yign in g cm-2; | ||||||
| # denotes the unit of in erg g-1. | |||||||||
Using the burst light curve data as listed in Table 1, we have estimated the spectral parameters of the blackbody emission in Table 7. The bolometric flux for each burst were obtained by applying the flux command to our spectral model (Model: 7 Tbabs*(*edge*nthcomp+bbodyrad)) in XSPEC over the energy range 0.1 - 100 keV. Following galloway2008thermonuclear, a bolometric correction was applied since the persistent emission extends over a broader energy range ( 80 keV) than the burst emission ( 25 keV). As our observational data cover only 3–79 keV, we have used the XSPEC command dummyrsp that generates a response beyond the observed energy limit (arnaud2003x). The burst fluence is the measure of the total energy radiated during the burst, measured in g cm-2(GRBCAT). The burst fluence () obtained for the bolometric flux over the whole energy range ( keV) during the burst integration times (Table 1) are listed in Table 7. Among the bursts, B1, B2 and B5 show fluences of 1.464, 1.456 and 1.435 (in units of 10-7 erg cm-2), respectively. The burst fluence does not follow a monotonic trend-it increases twice (in units of 10-7 erg cm-2): from 1.456 (B2) to 1.572 (B3), and again from 0.385 (B4) to 1.435 (B5). Typically, type-I thermonuclear bursts exhibit a gradual decrease in burst fluence as the neutron star surface depletes its nuclear fuel after each burst.
The burst fuel composition can be inferred from Qnuc, which is the nuclear energy released per nucleon for material with solar abundances (galloway2008thermonuclear). For a neutron star with and radius , Qnuc (in units of 4.4 MeV nucleon-1) is given by , where is the ratio of the integrated persistent flux to the burst fluence. The value of is given as , where cbol is the bolometric correction parameter, is the flux during persistent emission before the burst and is the recurrence time between consecutive bursts. While burst typically spans 2.5-25 keV, a broader energy range (upto 100 keV) must be considered for persistent emission due to its high-energy cutoff ( keV). In galloway2008thermonuclear, spectra were modeled over 0.1-200 keV, and cbol was calculated as a ratio between flux in the 0.1-200 keV and 2.5-25 keV energy bands. The derived bolometric correction factor () for the RXTE observation of the source 4U 1323-62 was 1.670.05 (galloway2008thermonuclear). Next, the recurring times between bursts - B1-B2; B3-B4; and B5-B6 are estimated as 4.88 hr, 4.44 hr and 4.24 hr, respectively yielding an average of = hr. Earlier studies reported burst recurrence times of 5.30–5.43 hr with EXOSAT (1985), 3.05 hr with ASCA (1994), 2.40–2.57 hr with BeppoSAX (1997), 2.45–2.59 hr with RXTE (1997), 5.59-5.65 hr with Suzaku (2007) and 2.65-2.70 hr with AstroSat (2017) (1989ApJ...338.1024P; balucinskachurch1999xraystudydippinglow; barnard2001; balucinska2009neutral; Bhulla_2020). Persistent emission spectra were extracted in the 3–79 keV energy range from intervals preceding each burst, as described in Section 4.1. The persistent fluxes () were calculated using the flux command applied to the best-fit results of Model 6 (Tbabs*(edge*nthcomp+bbodyrad)). For the burst peak intervals (segment S2, corresponding to 90% of the peak count rate), the peak fluxes () were similarly derived using the flux command on the best-fit results of Model 7 (Tbabs(*edge*nthcomp+bbodyrad)). From Table 7, we note values for all bursts range between 54 and 74, except B4. According to 2021ASSL..461..209G, such values (40–100) are characteristic of bursts powered by mixed hydrogen/helium fuel.
distinct properties with , the shortest decay time (29.5 s), and the lowest fluence (0.385 erg cm-2). Such a high typically indicates either accretion-driven emission or incomplete fuel burning (10.1093/mnras/staa2749). In contrast, the subsequent bursts (B5 and B6) with 64–74 are consistent with mixed H/He-fueled events.
The large of B4 suggests enhanced accretion energy and reduced nuclear burning efficiency, as following galloway2008thermonuclear, the parameter can also be expressed as , where is the gravitational potential energy released during accretion and z represents redshift which corresponds to for a neutron star with mass and radius R=10 km. Occurring at the bottom of an intensity dip, B4’s low flux likely results from coronal scattering (1987A&A...178..137F). Its peak count rate (94 counts s-1) is 38% below the mean (152 counts s-1), and its peak flux is 55% below the average, though its pre-burst persistent flux is not the lowest. The rise time (2 s) is also significantly shorter than in other bursts ( s). In this regard, jonker1998persistent reported the detection of a secondary burst in two observations, as well as a burst occurring during one of the dips, in the 1997 RXTE observation of the source. As noted by albayati2021discovery, certain bursts can display multiple peaks, and such double-peaked profiles are sometimes interpreted as separate bursts (bhattacharyya2007unusual). Therefore, B4’s low peak and short rise time may indicate that it represents a double-peaked burst event. This interpretation is further supported by the fact that burst B3 shows the highest fluence (1.572 erg cm-2) and the largest total energy release (1.882 erg), suggesting significant fuel burning immediately before the weaker B4 event.
The burst fluence represents the total radiated energy during a burst, expressed as (Eburst) in units of erg, calculated using , where the distance to the source () is assumed to be 10 kpc (Bhulla_2020). Except B4, all bursts yield Eburst values on the order of erg. The nuclear energy released for material with solar abundance is given by (in units of 4.4 MeV nucleon-1), where, denotes the hydrogen mass fraction (cumming2003models). For a fuel composition containing 90% hydrogen ( 0.9), the corresponding nuclear energy release is 5.2 MeV/nucleon ( erg g-1). As given in Table 7, the value of (in units of erg g-1) attains a maximum value of 3.45 for B3 and remains above 2.5 for all bursts except B4. According to Galloway_2020, such values indicate thermonuclear burning of mixed H/He fuel through rapid proton capture (rp) process, characteristic of short bursts. The comparatively low of B4 suggests predominant helium burning into heavier nuclei.
1981ApJ...247..267F examined hydrogen and helium burning on accreting neutron stars, developing models for pressure and temperature conditions necessary for unstable ignition at varying fuel-layer depths. Following galloway2008thermonuclear, the ignition column depth () in units of 108g cm-2 for a neutron star of radius is given as
The computed ignition depths () are listed in Table 7 for bursts B1-B6 are approximately 0.63, 0.56, 0.58, 0.58, 0.63 and 0.60 g cm-2, respectively. The identical ignition depths for B3 and B4 imply a comparable accreted mass column at those times. As the reported values are in the same order, we infer an average ignition depth . According to bildsten1997thermonuclearburningrapidlyaccreting and cumming2000rotational, such ignition depths are consistent with short Type-I thermonuclear bursts powered by mixed hydrogen/helium fuel.
6 Data Availability
We used the archival data from the NASA’s HEASARC database.
7 Acknowledgements
We thank NuSTARDAS for ensuring proper extraction of usable data. For the analysis of archival data we deeply acknowledge HEASOFT and CALDB. A token of thanks is extended towards the non-NET fellowship program of Visva-Bharati University for funding the research. We thank IUCAA for its vast repository of knowledge that enlightens curious minds.